OpenStudy (anonymous):

a toy rocket is pumped full of air and released upward. Its height can be approximated using the relation H=-7t squared + 42t + 49 where h is the height above ground in metres and t is the time in seconds

OpenStudy (anonymous):

whats the question here singh?

OpenStudy (anonymous):

your gonna awnser it

OpenStudy (anonymous):

a toy rocket is pumped full of air and released upward. Its height can be approximated using the relation H=-7t squared + 42t + 49 where h is the height above ground in metres and t is the time in seconds

OpenStudy (anonymous):

H=-7t^2 + 42t + 49 H=0, When -7t^2 + 42t + 49=0, t=7sec and t=-1sec so after 7 sec the toy rocket is in the ground

OpenStudy (anonymous):

how long after the launch did the rocket hit the ground explain

OpenStudy (anonymous):

after t=7sec the rocket hits the ground

OpenStudy (anonymous):

what time has passed when the rocket has reached its highest point

OpenStudy (anonymous):

H=-7t^2 + 42t + 4=0, that is solving for t in sec

OpenStudy (anonymous):

may i know if you are in the calculus class or algebra?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

ya so whats the awnser

OpenStudy (anonymous):

what time has passed when the rocket has reached its highest point

OpenStudy (anonymous):

ok... from H=-7t^2 + 42t + 49, this is a quadratic equation and a=-17,b=42 and c=49... the velocity equation is V=2ax+b=2(-7)t+42=-14t+42 now at max height V=0=-14t+42 we get t=3sec that is when the rocket reached its max height and the max height H=-7(3)^2 +42(3)+49=112 metres answer

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

Explain what happens to the graph of the exponential relation 2^x to the right of the y axis?

OpenStudy (anonymous):

meant to say 2 to the power of x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

are u there